OpenStudy (anonymous):

How would you factor x^2 -2x+1?

OpenStudy (anonymous):

(x-1)^2

OpenStudy (anonymous):

x^2 -2x+1= (x-1)(x-1)=(x-1)^2

OpenStudy (anonymous):

So we find the sum of -2 and product of 1?

OpenStudy (anonymous):

yes something like that,, do you know the FOIL method?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok good,, you are in good direction,,,

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

wc and good luck...

OpenStudy (anonymous):

\[x ^{2} - 2x + 1 (using (a-b)^{2}=a ^{2}-2ab+b ^{2})) = (x-1)^{?}\]